Roberto and Juanita are 9.9 miles apart. If they walk towards one another they will meet in one hour. If they both walk the same direction, Roberto will overtake Juanita in 11 hours. What is the rate of each.
Answer by ankor@dixie-net.com(22740)

Roberto and Juanita are 9.9 miles apart.
If they walk towards one another they will meet in one hour.
If they both walk the same direction, Roberto will overtake Juanita in 11 hours.
What is the rate of each.
:
Let r = R's walking speed
Let j = J's walking speed
:
Write a distance equation for each scenario
Toward
1(r + j) = 9.9
Same
11(r - j) = 9.9
simplify, divide both sides by 11, add the 1st equation:
r - j = .9
r + j = 9.9
------------adding eliminates j, find r
2r = 10.8
r = 10.8/2
r = 5.4 mph is R's walking speed
then
5.4 + j = 9.9
j = 9.9 - 5.4
j = 4.5 mph is J's walking speed
:
:
You can confirm this by substitution in both original equations